OpenStudy (anonymous):

Elementary linear algebra question: If \( A = \begin{bmatrix} 1 & 2\\ 3 & 4 \end{bmatrix} \) then, find \( I+A+A^2+ \cdots \infty \)

OpenStudy (anonymous):

haha how wld u solve this?

OpenStudy (anonymous):

Through magic lol, lets bring Turing here.

OpenStudy (anonymous):

omg tghis cat is getting menervous. I hate its eyes

OpenStudy (turingtest):

Your definition of elementary is pretty bold fool

OpenStudy (anonymous):

foool is bold

OpenStudy (anonymous):

can anybody help me on some problems nobody is answering them... plz :))

OpenStudy (anonymous):

lol, taken from a high school level book that's elementary, isn't ? :)

OpenStudy (jamesj):

Useful fact \[ \frac{1}{1-\lambda} = 1 + \lambda + \lambda^2 + .... \] provided \( |lambda| < 1 \)

OpenStudy (turingtest):

eigenvalues?

OpenStudy (turingtest):

or you just mean the sum of the geometric series?

OpenStudy (anonymous):

Yeah, why else James use \( \lambda \)? He could have used \(x\)

OpenStudy (jamesj):

yes, if you can find a diagonal matrix D such that \[ A = P^{-1}DP \] then this sum, if it exists is \[ P^{-1}(I + D + D^2 + ....)P \] and those terms are easy to calculate.

OpenStudy (anonymous):

why is this in a hs textbook??????

OpenStudy (turingtest):

I'm really not that good at linear algebra, I don't know that theorem james put up, and I'm trying to put the pieces together

OpenStudy (anonymous):

Turing : http://en.wikipedia.org/wiki/Diagonalizable_matrix

OpenStudy (turingtest):

Yeah this rings a bell gotta totally relearn in though

OpenStudy (anonymous):

Sweet :)

OpenStudy (anonymous):

sweet????? that didnt fit in lol

OpenStudy (anonymous):

lol, sour then :)

OpenStudy (phi):

\[\frac{1}{6}\left[\begin{matrix}3 & -2 \\ -3 & 0\end{matrix}\right]\]

OpenStudy (turingtest):

is that the answer?

OpenStudy (turingtest):

I keep getting an ugly eigenvalue

OpenStudy (turingtest):

*valueS

OpenStudy (phi):

Somewhere in Lecture 23 of Strang's LA he said (I-At)^(-1)= I + At + (At)^2 +...

OpenStudy (anonymous):

phi is right!

OpenStudy (anonymous):

At ?

OpenStudy (anonymous):

transpose of A?

OpenStudy (phi):

t is a variable He started by defining e^(At) where A is a matrix then did a btw

OpenStudy (turingtest):

hmm, I'll have to do OCW linear algebra after multivariable

OpenStudy (phi):

He's very good.

OpenStudy (anonymous):

Turing never gets enough of math lol

OpenStudy (anonymous):

So here you are doing \( (I-A)^ {-1} \) ?

OpenStudy (jamesj):

Yes, phi's approach is shorter and more elementary. In short, he wins the prize for this question.

OpenStudy (anonymous):

@JamesJ: Clearly it is, here is however a elementary proof that I came up with, \[ I+A+A^2+ \cdots \infty = X (say) \] \[ \implies A+A^2+A^3 \cdots \infty = AX \] Substracting, \[ X(I-A) = I \implies X = (I-A)^{-1} \space \space (QED) \]